How to solve the algebra equation (-1,6);m=5
Answer by Mathtut(3670)

I assume you want the equation that has the slope of 5 and goes thru the point
(-1,6)
:
y-6=5(x+1)
:
y-6=5x+5
:
y=5x+11....in slope intercept form
:
-5x+y=11...in standard form